### v5

@@ Appendix A Proofs for Section 3
Proof of Lemma 1.
For the sake of completeness, we provide the proof although the first part of the Lemma is the same as Theorem T-1 of Heckman and Pinto (2018) (applied with $\kappa(Y):=Y$). Let
$E_{z}(t|C)\equiv\mathbb{E}(Y_{i}{\rm 1\kern-3.99994pt1}(T_{i}=t)|Z_{i}=z,i\in C).$
We start from the sum over all response groups:
$\bar{E}_{z}(t)=\sum_{C}E_{z}(t|C)\Pr(i\in C).$
First note that if group $C$ does not have treatment $t$ under instrument $z$, it should not figure in the sum. Now if $C=C_{R}$ with $R(z)=t$, we have
$\displaystyle E_{z}(t|C)$ $\displaystyle=\mathbb{E}(Y_{i}{\rm 1\kern-3.99994pt1}(T_{i}=t)|Z_{i}=z,i\in C)$
$\displaystyle=\mathbb{E}(Y_{i}(t)|Z_{i}=z,i\in C)$
$\displaystyle=\mathbb{E}(Y_{i}(t)|i\in C).$
The second part of the Lemma is just adding up. ∎
Proof of Proposition 5.
@@
(A.1) $\displaystyle\begin{split}\bar{E}_{0}(1)&=\mathbb{E}(Y_{i}(1)|i\in C(\{1\},1))\Pr(i\in C(\{1\},1))=\mathbb{E}(Y_{i}(1)|i\in C_{11})\Pr(C_{11})\\ \text{ for }t\neq 1,\;\bar{E}_{0}(t)&=\mathbb{E}(Y_{i}(t)|i\in C(\emptyset,t))\Pr(i\in C(\emptyset,t))\\ &\;\;\;+\mathbb{E}(Y_{i}(t)|i\in C(\{1\},t))\Pr(i\in C(\{1\},t))\\ &=\mathbb{E}(Y_{i}(t)|i\in C_{tt})\Pr(C_{tt})+\mathbb{E}(Y_{i}(t)|i\in C_{t1})\Pr(C_{t1})\\ \bar{E}_{1}(1)&=\sum_{\tau\in\mathcal{T}}\mathbb{E}(Y_{i}(1)|i\in C(\{1\},\tau))\Pr(i\in C(\{1\},\tau))\\ &=\sum_{\tau\in\mathcal{T}}\mathbb{E}(Y_{i}(1)|i\in C_{\tau 1})\Pr(C_{\tau 1})\\ \text{ for }t\neq 1,\;\bar{E}_{1}(t)&=\mathbb{E}(Y_{i}(t)|i\in C(\emptyset,t))\Pr(i\in C(\emptyset,t))=\mathbb{E}(Y_{i}(t)|i\in C_{tt})\Pr(C_{tt}).\end{split}$
Since Proposition 5 identifies all type probabilities, the first and fourth equations in (A.1) give directly $\mathbb{E}(Y_{i}(t)|i\in C_{tt})$ for all $t$. Then the second equation identifies $\mathbb{E}(Y_{i}(t)|i\in C_{t1})$ for $t\neq 1$.
@@
Combining these results with Proposition 5 and Lemma 1 yields the formula in the Proposition. The denominator
@@
Proof of Corollary 2.
Recall from (A.1) that when $T=3$,
$\bar{E}_{1}(1)-\bar{E}_{0}(1)=\mathbb{E}\left[Y_{i}(1)|i\in C_{01}\right]\Pr(i\in C_{01})+\mathbb{E}\left[Y_{i}(1)|i\in C_{21}\right]\Pr(i\in C_{21}).$
Hence, under (3.5) we have
$\displaystyle\mathbb{E}\left[Y_{i}(1)|i\in C_{01}\right]\left\{\Pr(i\in C_{01})+\Pr(i\in C_{21})\right\}$
$\displaystyle\leq\bar{E}_{1}(1)-\bar{E}_{0}(1)$
$\displaystyle\leq\mathbb{E}\left[Y_{i}(1)|i\in C_{21}\right]\left\{\Pr(i\in C_{01})+\Pr(i\in C_{21})\right\}.$
The first conclusion of the corollary follows immediately, as
$\Pr(i\in C_{01})+\Pr(i\in C_{21})=P(1|1)-P(1|0).$
The testable prediction is a direct consequence of this chain of inequalities. ∎
Proof of Proposition 7.
It is straightforward from Figure 4 and Table 4. ∎
Proof of Proposition 8.
By Lemma 1, we obtain
@@
Replacing the probabilities and conditional expectations with their values from Proposition 7 and Proposition 8, we obtain $\Pr(i\in C_{010}\operatorname*{{\bigcup}}C_{012})=P(0|0)-P(0|1)$ and
@@
Replacing the probabilities and conditional expectations with their values from Proposition 7 and Proposition 8, we obtain $\Pr(i\in C_{002}\operatorname*{{\bigcup}}C_{012})=P(0|0)-P(0|2)$ and
@@
Appendix B Positive Selection in Head Start
@@
$Y_{i}(t)=k_{t}+m_{i}p_{t}+\zeta_{it},$
where $m_{i}>0$
$p_{h}>p_{c}>p_{n}=0.$
These conditions imply that children with a larger $m_{i}$ benefit more from preschool, especially from Head Start; $m_{i}$ does not play a role if $i$ goes to neither type of preschool. The $\zeta_{it}$ shocks are zero mean and idiosyncratic; we suppose that each subject $i$ expects $\mathbb{E}_{i}(Y_{i}(t))=k_{t}+m_{i}p_{t}$. Preference shocks depend positively on expected grades:
$u_{it}=a_{i}+b_{i}\mathbb{E}_{i}(Y_{i}(t))+\varepsilon_{it}=a_{i}+b_{i}k_{t}+b_{i}m_{i}p_{t}+\varepsilon_{it},$
@@
We assume that $b_{i}$, $m_{i}$, and the random vectors $(\eta_{ic},\eta_{ih})$ and $(\zeta_{in},\zeta_{ic},\zeta_{ih})$ are mutually independent.
@@
Let $A(\eta_{ic},\eta_{in},b_{i})$ and $B(\eta_{ic},\eta_{in},b_{i})$ be random subsets of $\mathbb{R}$ such that
$\sup A(\eta_{ic},\eta_{in},b_{i})\leq\inf B(\eta_{ic},\eta_{in},b_{i})$
@@
$\mathbb{E}(Y_{i}(t)\;|\;m_{i}\in A(\eta_{ic},\eta_{in},b_{i}))\leq\mathbb{E}(Y_{i}(t)|m_{i}\in B(\eta_{ic},\eta_{in},b_{i})).$
Proof of Lemma 2.
Take $t\in\{c,h\}$. Since $\mathbb{E}(Y_{i}(t)|m_{i}=m)=k_{t}+mp_{t}$, it is an increasing function of $m$. Fix $(\eta_{ic},\eta_{ih},b_{i})$; obviously, the distribution of $m_{i}$ conditional on $m_{i}\in B(\eta_{ic},\eta_{in},b_{i})$ first-order stochastically dominates that of $m_{i}$ conditional on $m_{i}\in A(\eta_{ic},\eta_{in},b_{i})$. Therefore
$\mathbb{E}(Y_{i}(t)\;|\;m_{i}\in A(\eta_{ic},\eta_{in},b_{i}),\eta_{ic},\eta_{ih},b_{i})\leq\mathbb{E}(Y_{i}(t)\;|\;m_{i}\in B(\eta_{ic},\eta_{in},b_{i}),\eta_{ic},\eta_{ih},b_{i}).$
@@
B.1 The Binary Instrument Case
@@
00$\displaystyle$ 0 $\displaystyle\geq\max(\underline{U}_{c}+v_{ic},\underline{U}_{h}+v_{ih}),$
@@
Note that $v_{ic}\geq-\underline{U}_{c}$ in $C_{ch}$ and $v_{ic}\leq-\underline{U}_{c}$ in $C_{nh}$. Since $v_{ic}=b_{i}d_{c}+b_{i}m_{i}p_{c}+\eta_{ic}$ and $p_{c}b_{i}>0$, it follows that for given $(\eta_{ic},\eta_{ih},b_{i})$, $m_{i}\geq m_{j}$ for any $i\in C_{ch}$ and $j\in C_{nh}$. Therefore we can apply Lemma 2 with $t=h$ to obtain
@@
B.2 The Ternary Instrument Case
@@
$\mathbb{E}(Y_{i}(n)\;|\;\eta_{ic},\eta_{ih},b_{i})=k_{n}$
is constant. Therefore, trivially,
$\mathbb{E}(Y_{i}(n)|i\in\tilde{C}_{nhc})\geq\mathbb{E}(Y_{i}(n)|i\in\tilde{C}_{nnn}).$
Now let us consider the response-groups $\tilde{C}_{n\ast c}$ and $\tilde{C}_{chc}$. $\tilde{C}_{n\ast c}$ is defined by the inequalities
(B.1) $\bar{U}_{c}+v_{ic}>0>\max(\underline{U}_{h}+v_{ih},\underline{U}_{c}+v_{ic});$
@@
(B.2) $\bar{U}_{h}+v_{ih}>\underline{U}_{c}+v_{ic}>\max(\underline{U}_{h}+v_{ih},0).$
(B.1) implies that $v_{ic}=b_{i}d_{c}+b_{i}m_{i}p_{c}+\eta_{ic}<-\underline{U}_{c}$, while (B.2) implies the reverse inequality. Here also, applying Lemma 2 with $t=c$ directly gives the conclusion:
$\mathbb{E}(Y_{i}(c)|i\in\tilde{C}_{n\ast c})\leq\mathbb{E}(Y_{i}(n)|i\in\tilde{C}_{chc}).$
Appendix C Proofs for Section 3.4
@@
It therefore excludes the response-groups $C_{02\ast},C_{20\ast},C_{0\ast 1}$, and $C_{1\ast 0}$. The response-group $C_{021}$ appears twice in this list; and four other response-groups were already ruled out by Assumption 4. The reader can easily check that the $8\displaystyle 3^{3}-12-(11-4)=$ 8 response-groups left are exactly the same as in our Figure 4.
Proof of Proposition 9.
@@
(C.1) $\mathbb{E}\left[\left(Y_{i}-\beta_{0}-\beta_{1}{\rm 1\kern-3.99994pt1}(T_{i}=1)-\beta_{2}{\rm 1\kern-3.99994pt1}(T_{i}=2)\right){\rm 1\kern-3.99994pt1}(Z_{i}=z)\right]=0$
@@
(C.2) $Y_{i}=Y_{i}(0)+(Y_{i}(1)-Y_{i}(0)){\rm 1\kern-3.99994pt1}(T_{i}=1)+(Y_{i}(1)-Y_{i}(0)){\rm 1\kern-3.99994pt1}(T_{i}=2),$
which allows us to write Equation C.1 as
(C.3) $\mathbb{E}\left[\left(Y_{i}(0)-\beta_{0}+b_{i}(1){\rm 1\kern-3.99994pt1}(T_{i}=1)+b_{i}(2){\rm 1\kern-3.99994pt1}(T_{i}=2)\right){\rm 1\kern-3.99994pt1}(Z_{i}=z)\right]=0,$
@@
Proof of Corollary 4.
Solving the system of equations in Proposition 9 gives, after elementary calculations,
@@
(C.4) $\displaystyle\begin{split}&\beta_{1}\mathcal{D}\\ &\geq\Pr(i\in\mathcal{C}_{1})\Pr(i\in\mathcal{C}_{2})\mathbb{E}\left(Y_{i}(1)-Y_{i}(0)|i\in\mathcal{C}_{1}\right)-\Pr(i\in C_{112})\Pr(i\in C_{212})\mathbb{E}\left(Y_{i}(1)-Y_{i}(0)|i\in C_{112}\right).\end{split}$
@@
Note that $a$ and $b$ are non-negative, and $a-b=\mathcal{D}\neq 0$. Suppose that $\mathcal{D}>0$ so that Equation C.4 becomes $\beta_{1}\geq r$. Since $\alpha-\beta=\mathcal{D}_{1}\geq 0$, we can apply result 1 and we get
@@
Appendix D Revisiting the $2\times 3$ and $3\times 3$ Models via Heckman and Pinto (2018)
D.1 Notation

### v6

@@ Appendix B Proofs for Section˜3
Proof of Proposition˜5.
@@
(B.1) $\displaystyle\begin{split}\bar{E}_{0}(1)&=\mathbb{E}(Y_{i}(1)|i\in C(\{1\},1))\Pr(i\in C(\{1\},1))=\mathbb{E}(Y_{i}(1)|i\in C_{11})\Pr(C_{11})\\ \text{ for }t\neq 1,\;\bar{E}_{0}(t)&=\mathbb{E}(Y_{i}(t)|i\in C(\emptyset,t))\Pr(i\in C(\emptyset,t))\\ &\;\;\;+\mathbb{E}(Y_{i}(t)|i\in C(\{1\},t))\Pr(i\in C(\{1\},t))\\ &=\mathbb{E}(Y_{i}(t)|i\in C_{tt})\Pr(C_{tt})+\mathbb{E}(Y_{i}(t)|i\in C_{t1})\Pr(C_{t1})\\ \bar{E}_{1}(1)&=\sum_{\tau\in\mathcal{T}}\mathbb{E}(Y_{i}(1)|i\in C(\{1\},\tau))\Pr(i\in C(\{1\},\tau))\\ &=\sum_{\tau\in\mathcal{T}}\mathbb{E}(Y_{i}(1)|i\in C_{\tau 1})\Pr(C_{\tau 1})\\ \text{ for }t\neq 1,\;\bar{E}_{1}(t)&=\mathbb{E}(Y_{i}(t)|i\in C(\emptyset,t))\Pr(i\in C(\emptyset,t))=\mathbb{E}(Y_{i}(t)|i\in C_{tt})\Pr(C_{tt}).\end{split}$
Since Proposition˜5 identifies all type probabilities, the first and fourth equations in (B.1) give directly $\mathbb{E}(Y_{i}(t)|i\in C_{tt})$ for all $t$. Then the second equation identifies $\mathbb{E}(Y_{i}(t)|i\in C_{t1})$ for $t\neq 1$.
@@
Combining these results with Proposition˜5 and Lemma˜1 yields the formula in the Proposition. The denominator
@@
Proof of Proposition˜7.
First, we consider the baseline counterfactuals. By Proposition˜6, the expectations of $Y_{i}(0)$ and $Y_{i}(2)$ for the complier groups are point identified as:
$\displaystyle\mathbb{E}[Y_{i}(0)\mid i\in C_{01}]$ $\displaystyle=\frac{\bar{E}_{0}(0)-\bar{E}_{1}(0)}{P(0\mid 0)-P(0\mid 1)},$
$\displaystyle\mathbb{E}[Y_{i}(2)\mid i\in C_{21}]$ $\displaystyle=\frac{\bar{E}_{0}(2)-\bar{E}_{1}(2)}{P(2\mid 0)-P(2\mid 1)}.$
Next, we consider the treated counterfactuals. Recall from (B.1) that when $T=3$,
$\bar{E}_{1}(1)-\bar{E}_{0}(1)=\mathbb{E}[Y_{i}(1)\mid i\in C_{01}]\Pr(i\in C_{01})+\mathbb{E}[Y_{i}(1)\mid i\in C_{21}]\Pr(i\in C_{21}).$
This implies that the weighted sum of $L_{01}$ and $L_{21}$ satisfies the equality given in (3.7). Furthermore, under assumption (3.6), we have
$\displaystyle\mathbb{E}[Y_{i}(1)\mid i\in C_{01}]\{\Pr(i\in C_{01})+\Pr(i\in C_{21})\}$
$\displaystyle\quad\leq\bar{E}_{1}(1)-\bar{E}_{0}(1)$
$\displaystyle\quad\leq\mathbb{E}[Y_{i}(1)\mid i\in C_{21}]\{\Pr(i\in C_{01})+\Pr(i\in C_{21})\}.$
Substituting the point-identified expressions for $\mathbb{E}[Y_{i}(0)\mid i\in C_{01}]$ and $\mathbb{E}[Y_{i}(2)\mid i\in C_{21}]$ derived above, and using the identity $\Pr(i\in C_{01})+\Pr(i\in C_{21})=P(1\mid 1)-P(1\mid 0)$, yields the inequalities in (3.8). The point identification results under the special case when (3.6) holds with equality follow immediately. ∎
Proof of Proposition˜8.
It is straightforward from Figure˜4 and Table˜5. ∎
Proof of Proposition˜9.
By Lemma˜1, we obtain
@@
Replacing the probabilities and conditional expectations with their values from Proposition˜8 and Proposition˜9, we obtain $\Pr(i\in C_{010}\operatorname*{{\bigcup}}C_{012})=P(0|0)-P(0|1)$ and
@@
Replacing the probabilities and conditional expectations with their values from Proposition˜8 and Proposition˜9, we obtain $\Pr(i\in C_{002}\operatorname*{{\bigcup}}C_{012})=P(0|0)-P(0|2)$ and
@@
Appendix C Positive Selection in Head Start
@@
$Y_{i}(t)=f_{i}+k_{t}+m_{i}p_{t}+\zeta_{it},$
where $m_{i}>0$, $p_{h}>0$, $p_{c}>0$, and $p_{n}=0$. These conditions imply that children with a larger $m_{i}$ benefit more from preschool (either Head Start or other centers); however, $m_{i}$ does not play a role if $i$ goes to neither type of preschool. The $\zeta_{it}$ shocks are zero mean and idiosyncratic; we suppose that each subject $i$ expects $\mathbb{E}_{i}(Y_{i}(t))=f_{i}+k_{t}+m_{i}p_{t}$. Preference shocks depend positively on expected grades:
$u_{it}=a_{i}+b_{i}\mathbb{E}_{i}(Y_{i}(t))+\varepsilon_{it}=a_{i}+b_{i}f_{i}+b_{i}k_{t}+b_{i}m_{i}p_{t}+\varepsilon_{it},$
@@
We assume that $b_{i}$, $m_{i}$, and the random vectors $(\eta_{ic},\eta_{ih})$ and $(\zeta_{in},\zeta_{ic},\zeta_{ih})$ are mutually independent; and that $f_{i}$ is independent of $(\eta_{ic},\eta_{ih},\zeta_{in},\zeta_{ic},\zeta_{ih})$ while $\mathbb{E}(f_{i}\mid m_{i}=m)$ is non-decreasing in $m$.
@@
Let $A(\eta_{ic},\eta_{ih},b_{i})$ and $B(\eta_{ic},\eta_{ih},b_{i})$ be random subsets of $\mathbb{R}$ such that
$\sup A(\eta_{ic},\eta_{ih},b_{i})\leq\inf B(\eta_{ic},\eta_{ih},b_{i})$
@@
$\mathbb{E}(Y_{i}(t)\mid m_{i}\in A(\eta_{ic},\eta_{ih},b_{i}))\leq\mathbb{E}(Y_{i}(t)\mid m_{i}\in B(\eta_{ic},\eta_{ih},b_{i})).$
Proof of Lemma˜2.
Take $t\in\{c,h\}$. Since $\mathbb{E}(Y_{i}(t)\mid m_{i}=m)=\mathbb{E}(f_{i}\mid m_{i}=m)+k_{t}+mp_{t}$, it is a non-decreasing function of $m$ (recalling that $\mathbb{E}(f_{i}\mid m_{i}=m)$ is non-decreasing and $p_{t}>0$). Fix $(\eta_{ic},\eta_{ih},b_{i})$; obviously, the distribution of $m_{i}$ conditional on $m_{i}\in B(\eta_{ic},\eta_{ih},b_{i})$ first-order stochastically dominates that of $m_{i}$ conditional on $m_{i}\in A(\eta_{ic},\eta_{ih},b_{i})$. Therefore,
$\mathbb{E}(Y_{i}(t)\mid m_{i}\in A(\eta_{ic},\eta_{ih},b_{i}),\eta_{ic},\eta_{ih},b_{i})\leq\mathbb{E}(Y_{i}(t)\mid m_{i}\in B(\eta_{ic},\eta_{ih},b_{i}),\eta_{ic},\eta_{ih},b_{i}).$
@@
C.1 The Binary Instrument Case
@@
$\displaystyle 0$ $\displaystyle\geq\max(\underline{U}_{c}+v_{ic},\underline{U}_{h}+v_{ih}),$
@@
Note that $v_{ic}\geq-\underline{U}_{c}$ in $C_{ch}$ and $v_{ic}\leq-\underline{U}_{c}$ in $C_{nh}$. Fix values $\bar{\eta}_{c},\bar{\eta}_{h},$ and $\bar{b}$ and take $i\in C_{ch}$ and $j\in C_{nh}$ such that
$\eta_{ic}=\eta_{jc}=\bar{\eta}_{c};\;\eta_{ih}=\eta_{jh}=\bar{\eta}_{h};\;b_{i}=b_{j}=\bar{b}.$
Since $v_{kc}=\bar{b}d_{c}+\bar{b}m_{k}p_{c}+\bar{\eta}_{c}$ for $k=i,j$ and $p_{c}\bar{b}>0$, it follows that $m_{i}\geq m_{j}$. Therefore we can apply Lemma 2 with $t=h$ to obtain
@@
C.2 The Ternary Instrument Case
@@
$\mathbb{E}(Y_{i}(n)\mid\eta_{ic},\eta_{ih},b_{i})=\mathbb{E}(f_{i}\mid\eta_{ic},\eta_{ih},b_{i})+k_{n}.$
Given our assumptions on the distribution of $f_{i}$, we have
$\mathbb{E}(f_{i}\mid\eta_{ic},\eta_{ih},b_{i})=\mathbb{E}\left[\mathbb{E}(f_{i}\mid\eta_{ic},\eta_{ih},b_{i},m_{i})\mid\eta_{ic},\eta_{ih},b_{i}\right]=\mathbb{E}\left[\mathbb{E}(f_{i}\mid m_{i})\mid\eta_{ic},\eta_{ih},b_{i}\right].$
Figure 7 shows that if $i\in\tilde{C}_{nnn}$ and $j\in\tilde{C}_{nhc}$, then $v_{ih}\leq v_{jh}$ and $v_{ic}\leq v_{jc}$. Since $v_{it}=b_{i}d_{t}+b_{i}m_{i}p_{t}+\eta_{it},$ the distribution of $m$ in $\tilde{C}_{nnn}$ is dominated by its distribution in $\tilde{C}_{nhc}$. Since $\mathbb{E}(f_{i}\mid m_{i}=m)$ is non-decreasing in $m$, it follows that
$\mathbb{E}(Y_{i}(n)\mid i\in\tilde{C}_{nhc})\geq\mathbb{E}(Y_{i}(n)\mid i\in\tilde{C}_{nnn}).$
Now let us consider the response-groups $\tilde{C}_{n*c}$ and $\tilde{C}_{chc}$. $\tilde{C}_{n*c}$ is defined by the inequalities
(C.1) $\bar{U}_{c}+v_{ic}>0>\max(\underline{U}_{h}+v_{ih},\underline{U}_{c}+v_{ic});$
@@
(C.2) $\bar{U}_{h}+v_{ih}>\underline{U}_{c}+v_{ic}>\max(\underline{U}_{h}+v_{ih},0).$
(C.1) implies that $v_{ic}=b_{i}d_{c}+b_{i}m_{i}p_{c}+\eta_{ic}<-\underline{U}_{c}$, while (C.2) implies the reverse inequality. Here also, applying Lemma 2 with $t=c$ directly gives the conclusion:
$\mathbb{E}(Y_{i}(c)\mid i\in\tilde{C}_{n*c})\leq\mathbb{E}(Y_{i}(c)\mid i\in\tilde{C}_{chc}).$
Appendix D Proofs for Section˜3.4
@@
It therefore excludes the response-groups $C_{02\ast},C_{20\ast},C_{0\ast 1}$, and $C_{1\ast 0}$. The response-group $C_{021}$ appears twice in this list; and four other response-groups were already ruled out by Assumption 4. The reader can easily check that the $3^{3}-12-(11-4)=8$ response-groups left are exactly the same as in our Figure˜4.
Proof of Proposition˜11.
@@
(D.1) $\mathbb{E}\left[\left(Y_{i}-\beta_{0}-\beta_{1}{\rm 1\kern-3.99994pt1}(T_{i}=1)-\beta_{2}{\rm 1\kern-3.99994pt1}(T_{i}=2)\right){\rm 1\kern-3.99994pt1}(Z_{i}=z)\right]=0$
@@
(D.2) $Y_{i}=Y_{i}(0)+(Y_{i}(1)-Y_{i}(0)){\rm 1\kern-3.99994pt1}(T_{i}=1)+(Y_{i}(1)-Y_{i}(0)){\rm 1\kern-3.99994pt1}(T_{i}=2),$
which allows us to write Equation˜D.1 as
(D.3) $\mathbb{E}\left[\left(Y_{i}(0)-\beta_{0}+b_{i}(1){\rm 1\kern-3.99994pt1}(T_{i}=1)+b_{i}(2){\rm 1\kern-3.99994pt1}(T_{i}=2)\right){\rm 1\kern-3.99994pt1}(Z_{i}=z)\right]=0,$
@@
Proof of Proposition˜12.
Solving the system of equations in Proposition˜11 gives, after elementary calculations,
@@
(D.4) $\displaystyle\begin{split}&\beta_{1}\mathcal{D}\\ &\geq\Pr(i\in\mathcal{C}_{1})\Pr(i\in\mathcal{C}_{2})\mathbb{E}\left(Y_{i}(1)-Y_{i}(0)|i\in\mathcal{C}_{1}\right)-\Pr(i\in C_{112})\Pr(i\in C_{212})\mathbb{E}\left(Y_{i}(1)-Y_{i}(0)|i\in C_{112}\right).\end{split}$
@@
Note that $a$ and $b$ are non-negative, and $a-b=\mathcal{D}\neq 0$. Suppose that $\mathcal{D}>0$ so that Equation˜D.4 becomes $\beta_{1}\geq r$. Since $\alpha-\beta=\mathcal{D}_{1}\geq 0$, we can apply result 1 and we get
@@
Appendix E Revisiting the $2\times 3$ and $3\times 3$ Models via Heckman and Pinto (2018)
E.1 Notation